# Non-Convex Self-Concordant Functions: Practical Algorithms and Complexity Analysis

### v1

@@ A.3 Convergence Analysis
In this section, we prove the convergence of Algorithms 3 and 4. To do so, we first present some technical lemmas. Lemma A.4 analyzes the function sqrt-cond in Algorithm 3 by a classical result [Kuczynski-1992-Estimating] on the Lanczos method. Lemma A.5 analyzes the function CG-Inverse in Algorithm 3 using [Wright-2006-Numerical, Theorem 5.5]. Lemma A.6 describes the relationship between the condition numbers of the Hessians at two nearby points for self-concordant functions.
@@
Let $\epsilon^{\prime}=\frac{1}{10\beta^{2}}$. It follows from [Kuczynski-1992-Estimating, Theorem 4.2] that, with probability at least $\frac{1}{2}\delta$, the Lanczos method outputs $\lambda_{1}$ such that $|\lambda_{1}-\lambda_{\max}(A)|\leq\frac{1}{3}\lambda_{\max}(A)$ in $O(\min\{n,\,\log(n\delta^{-2})\})$ matrix-vector multiplications. With probability at least $1-\frac{1}{2}\delta$, the Lanczos method outputs $\lambda^{\prime}$ such that $\left|\lambda^{\prime}-\lambda_{\max}\left(\tfrac{2\lambda_{1}}{1-\epsilon^{\prime}}I-A\right)\right|\leq\epsilon^{\prime}\cdot\allowbreak\lambda_{\max}\left(\tfrac{2\lambda_{1}}{1-\epsilon^{\prime}}I-A\right)\leq\tfrac{1}{3}\lambda_{\min}(A)$ in $O(\min\{n,\beta\log(n\delta^{-2})\})$ matrix-vector multiplications. Letting $\lambda_{2}=\frac{2\lambda_{2}}{1-\epsilon^{\prime}}-\lambda^{\prime}$, then we have $|\lambda_{2}-\lambda_{\min}(A)|\leq\frac{1}{3}\lambda_{\min}(A)$. Thus, letting $X^{2}=2\lambda_{1}/\lambda_{2}$ suffices. ∎
Lemma A.5 (Theorem 5.5 of [Wright-2006-Numerical]).
@@
[Nesterov-2018-Lectures, Theorem 5.1.13] implies that $x_{f}^{\star}$ exists. It follows from [Nesterov-2018-Lectures, Theorem 5.1.7, Theorem 5.2.1] that $(1-\kappa r)^{2}\nabla^{2}f(x)\preceq\nabla^{2}f(x_{f}^{\star})\preceq\frac{1}{(1-\kappa r)^{2}}\nabla^{2}f(x)$, where $r=\|x-x_{f}^{\star}\|_{\nabla^{2}f(x)}\leq\frac{1}{(R^{-1}-1)\kappa}$. The desired result follows from direct calculation. ∎
@@
Suppose $f$ is $\kappa$-self-concordant. At any given iteration of Algorithm 3, if $\beta_{k}^{2}\geq\mathrm{cond}(\nabla^{2}f(x_{k}))$ and $\rho_{k}>\frac{R_{1}^{2}}{\kappa^{2}}$, then
$f(x_{k})-f(x_{k+1})\geq\kappa^{-2}\,\omega\left(\tfrac{\sqrt{1-\alpha_{\star}}}{1+\alpha_{\star}}\,\kappa\sqrt{\rho_{k}}\right)\geq\kappa^{-2}\omega(R_{3})$
and $\sqrt{\tfrac{\mathrm{cond}(\nabla^{2}f(x_{k+1}))}{\mathrm{cond}(\nabla^{2}f(x_{k}))}}\leq B_{k}\leq\exp\left(C_{1}\kappa^{2}(f(x_{k})-f(x_{k+1}))\right)$.
@@
It follows from Lemma A.5 that ${\frac{\rho_{k}}{\delta_{k}}}\in[\frac{1-\alpha_{\star}}{(1+\alpha_{\star})^{2}},\frac{1+\alpha_{\star}}{(1-\alpha_{\star})^{2}}]$. Thus, Proposition 2.6 implies that $f(x_{k})-f(x_{k+1})\geq\kappa^{-2}\omega\left(\tfrac{\kappa\rho_{k}}{\sqrt{\delta_{k}}}\right)\geq\kappa^{-2}\omega\left(\tfrac{\sqrt{1-\alpha_{\star}}}{1+\alpha_{\star}}\kappa\sqrt{\rho_{k}}\right)\geq\kappa^{-2}\omega(R_{3})$. Let $r_{k}=\|x_{k+1}-x_{k}\|_{\nabla^{2}f(x_{k})}=\frac{\rho_{k}\sqrt{\delta_{k}}}{\delta_{k}+\kappa\rho_{k}\sqrt{\delta_{k}}}$. Hence, we have $(1-\kappa r_{k})^{-1}\leq 1+\frac{\sqrt{1+\alpha_{\star}}}{1-\alpha_{\star}}\kappa\sqrt{\rho_{k}}$. It follows from [Nesterov-2018-Lectures, Theorem 5.1.7] that $\sqrt{\tfrac{\mathrm{cond}(\nabla^{2}f(x_{k+1}))}{\mathrm{cond}(\nabla^{2}f(x_{k}))}}\leq(1-\kappa r_{k})^{-2}\leq B_{k}$. Moreover, we have
$\tfrac{\log B_{k}}{\kappa^{2}(f(x_{k})-f(x_{k+1}))}\leq\tfrac{2\log\left(1+\frac{\sqrt{1+\alpha_{\star}}}{1-\alpha_{\star}}\,\kappa\sqrt{\rho_{k}}\right)}{\omega\left(\tfrac{\sqrt{1-\alpha_{\star}}}{1+\alpha_{\star}}\kappa\sqrt{\rho_{k}}\right)}\leq\tfrac{2\log\left(1+\frac{\sqrt{1+\alpha_{\star}}}{1-\alpha_{\star}}R_{1}\right)}{\omega\left(\tfrac{\sqrt{1-\alpha_{\star}}}{1+\alpha_{\star}}R_{1}\right)}\leq C_{1},$
where the second inequality holds by the monotonicity of $\frac{\log\left(1+\frac{\sqrt{1+\alpha_{\star}}}{1-\alpha_{\star}}t\right)}{\omega\left(\frac{\sqrt{1-\alpha_{\star}}}{1+\alpha_{\star}}t\right)}$ for $t>R_{1}$. This completes the proof. ∎
@@
If $\rho_{k}\leq\frac{R_{1}^{2}}{\kappa^{2}}$, then we have $t_{k}=\frac{\rho_{k}}{\delta_{K}+2\kappa\rho_{k}\sqrt{\delta_{k}}}$. Letting $\lambda_{k}=\lambda_{f}(x_{k})$, $H_{k}=\nabla^{2}f(x_{k})$, $G_{k}=\int_{0}^{1}\nabla^{2}f(x_{k}+\tau(x_{k+1}-x_{k}))d\tau$, $\zeta_{k}=\kappa\sqrt{\delta_{k}}$, we have $\nabla f(x_{k+1})=G_{k}(x_{k+1}-x_{k})+\nabla f(x_{k})=-t_{k}G_{k}h_{k}+g_{k}$. It follows from [Nesterov-2018-Lectures, Corollary 5.1.5] that $(1-t_{k}\zeta_{k}+\frac{1}{3}t_{k}^{2}\zeta_{k}^{2})H_{k}\preceq G_{k}\preceq\frac{1}{1-t_{k}\zeta_{k}}H_{k}$. Moreover, [Nesterov-2018-Lectures, Lemma 5.1.7] implies that $H_{k+1}^{-1}\preceq\frac{1}{(1-t_{k}\zeta_{k})^{2}}H_{k}^{-1}$. Thus, we have $\lambda_{f}(x_{k+1})^{2}\leq\tfrac{1}{(1-t_{k}\zeta_{k})^{2}}\|\nabla f(x_{k+1})\|_{H_{k}^{-1}}^{2}$, where
$\|\nabla f(x_{k+1})\|_{H_{k}^{-1}}^{2}=\lambda_{k}^{2}-2t_{k}h_{k}^{\top}G_{k}h_{k}+t_{k}^{2}h_{k}^{\top}G_{k}H_{k}^{-1}G_{k}h_{k}-2t_{k}h_{k}^{\top}G_{k}(H_{k}^{-1}g_{k}-h_{k}).$
@@
$|2h_{k}^{\top}G_{k}(H_{k}^{-1}g_{k}-h_{k})|\leq\alpha_{\star}\|G_{k}h_{k}\|_{H_{k}^{-1}}^{2}+\alpha_{\star}^{-1}\|H_{k}^{-1}g_{k}-h_{k}\|_{H_{k}}^{2}.$
@@
Lemma A.5 implies $\zeta_{k}\in[0,\frac{1+\alpha_{\star}}{\sqrt{1-\alpha_{\star}}}R_{1}]$. For any $\zeta_{l}<\zeta_{u}$, we define $t_{u}=\frac{1}{\frac{(1-\alpha_{\star})^{2}}{1+\alpha_{\star}}+3\zeta_{l}}$ and $t_{l}=\frac{1}{\frac{(1+\alpha_{\star})^{2}}{1-\alpha_{\star}}+3\zeta_{u}}$. Since $t_{k}=\frac{1}{\frac{\delta_{k}}{\rho_{k}}+3\zeta_{k}}$ and $\frac{\delta_{k}}{\rho_{k}}\in[\frac{(1-\alpha_{\star})^{2}}{1+\alpha_{\star}},\frac{(1+\alpha_{\star})^{2}}{1-\alpha_{\star}}]$, we have
$g(t_{k},\zeta_{k})\leq\tfrac{1-2t_{l}(1-\alpha_{\star})^{2}(1-t_{u}\zeta_{u})}{(1-t_{l}\zeta_{l})^{2}}+\tfrac{t_{u}^{2}(1+\alpha_{\star})^{2}}{(1-t_{u}\zeta_{u})^{4}}+\tfrac{\alpha_{\star}t_{u}}{{(1-t_{u}\zeta_{u})^{2}}}\left(\tfrac{(1+\alpha_{\star})^{2}}{(1-t_{u}\zeta_{u})^{2}}+1\right)$
for any $\zeta_{k}\in[\zeta_{l},\zeta_{u}]$. Dividing $0,\frac{1+\alpha_{\star}}{\sqrt{1-\alpha_{\star}}}R_{1}]$ into $1000$ intervals equally, we can verify numerically that $g(t_{k},\zeta_{k})\leq 0.945$. The desired result follows from $\lambda_{k}^{2}\leq\rho_{k}(1-\alpha_{\star})^{-1}$. ∎
@@
Lemma A.9 shows that $A_{1}\cap A_{2}\cap A_{3}$ happens with high probability. Hence, using two calls of sqrt-cond, Algorithm 3 gets good estimates of Hessian condition numbers.
If $A_{1}\cap A_{2}\cap A_{3}$ holds, then for all $0\leq k\leq k^{\star}$, using Lemma A.7, we have
@@
By Lemma A.4, we have $\mathbb{P}\left(A_{1}\right)\geq 1-\frac{1}{2}\delta$. Therefore, it suffices to show $\mathbb{P}\left(A_{2}|A_{1}\right)\geq 1-\frac{1}{2}\delta$ and $A_{1}\cap A_{2}\subseteq A_{3}$. To show $\mathbb{P}\left(A_{2}|A_{1}\right)\geq 1-\frac{1}{2}\delta$, we observe from Lemma A.7 that $\beta_{0}^{2}\geq\mathrm{cond}(\nabla^{2}f(x_{0}))$ implies $\beta_{k^{\star}}^{2}\geq\mathrm{cond}(\nabla^{2}f(x_{k^{\star}}))$. Thus, Lemma A.4 implies $\mathbb{P}\left(A_{2}|A_{1}\right)\geq 1-\frac{1}{2}\delta$. It remains to show $A_{1}\cap A_{2}\subseteq A_{3}$. Suppose that $A_{1}\cap A_{2}$ holds, then we have $\mathrm{cond}(\nabla^{2}f(x_{l}))\leq\beta_{l}^{2}$ for $0\leq l\leq k^{\star}$ by Lemma A.7. Moreover, Lemma A.8 yields $\rho_{l}\leq\frac{R_{1}^{2}}{\kappa^{2}}$ for all $l\geq k^{\star}+1$. Therefore, using Lemma A.6, we have $\mathrm{cond}(\nabla^{2}f(x_{l}))\leq C_{3}^{8}\mathrm{cond}(\nabla^{2}f(x_{k^{\star}}))\leq C_{3}^{8}(\beta^{\star})^{2}=\beta_{\mathrm{local}}^{2}=\beta_{l}^{2}$. This completes the proof. ∎
@@
Suppose that $f$ is $\kappa$-self-concordant, bounded below, and that $\beta^{2}\geq\mathrm{cond}(\nabla^{2}f(x_{0}))$. Then, with probability at least $1-\delta$, the following statements hold:
  1. 1.
    Algorithm 3 terminates and returns a point $y$. If $y\neq x_{0}$, then $\lambda_{f}(y)\leq\epsilon_{1}$ and $\lambda_{f}(x_{0})\geq\sqrt{\frac{1-\alpha_{\star}}{1+\alpha_{\star}}}\,\epsilon_{0}$; if $y=x_{0}$, then $\lambda_{f}(x_{0})\leq\epsilon_{0}$. Let $\Delta_{f}=\kappa^{2}(f(x_{0})-f(y))$, $K_{1}=\frac{\Delta_{f}}{\omega(R_{3})}$, and $K_{2}=2\log_{C_{2}^{-1}}\!\left(\frac{R_{2}}{\kappa\epsilon_{1}}\right)$. Then Algorithm 3 terminates within $2+\lfloor K_{1}\rfloor+\max(\lfloor K_{2}\rfloor,0)$ iterations.
  2. 2.
    Define $B=\max\left(\sqrt{\mathrm{cond}(\nabla^{2}f(x_{0}))},\sqrt{\mathrm{cond}(\nabla^{2}f(\arg\min_{y}f(y)))}\right)$, and
    $\Lambda_{1}$ $=C_{\mathrm{CG}}\min\left\{n,\,2Be^{C_{1}\Delta_{f}}\log(\alpha_{\star}^{-1})\right\},$
    $\Lambda_{2}$ $=C_{\mathrm{CG}}\min\left\{n,\,2BC_{3}^{6}\log(\alpha_{\star}^{-1})\right\},$
    $\Lambda_{3}$ $=C_{L}\left(\min\left\{n,\,\beta\log(n\delta^{-2})\right\}+\min\left\{n,\,2Be^{C_{1}\Delta_{f}}\log(n\delta^{-2})\right\}\right).$
    The total number of HVP computations performed by Algorithm 3 is at most $(\Lambda_{1}+1)(K_{1}+1)+(\Lambda_{2}+1)(K_{2}+1)+\Lambda_{3}=O(\log(\epsilon_{1}^{-1})+\log(\delta^{-1}))$.
@@
By Lemma A.9, it suffices to prove all statements hold deterministically assuming that the event $A_{1}\cap A_{2}\cap A_{3}$ holds. For all $k=0,\ldots,k^{\star}-1$, we have $\rho_{k}>\frac{R_{1}^{2}}{\kappa^{2}}$. Therefore, it follows from Lemma A.7 that $f(x_{k})-f(x_{k+1})\geq\kappa^{-2}\omega(R_{3})$. Summing this inequality over $k=0,1,\dots,k^{\star}-1$ yields ${k^{\star}}\kappa^{-2}\omega(R_{3})\leq f(x_{0})-f(x_{k^{\star}})$. Thus, $k^{\star}$ must be finite. Lemma A.8 implies that we have $\rho_{k+1}\leq C_{2}\rho_{k}\leq\frac{R_{1}^{2}}{\kappa^{2}}$ for all $k\geq k^{\star}$. Thus, we have $(1-\alpha_{\star})\epsilon_{1}^{2}\leq\rho_{k}\leq C_{2}^{k-k^{\star}}\rho_{k^{\star}}\leq C_{2}^{k-k^{\star}}\frac{R_{1}^{2}}{\kappa^{2}}$. Therefore, the algorithm terminates at some $k\leq 1+\lfloor K_{1}\rfloor+\max(\lfloor K_{2}\rfloor,0)$.
@@
Now, we can estimate the total number of HVPs. By Lemma A.5, the total number of HVPs called by CG-Inverse is bounded by $\sum_{k=0}^{k^{\star}}C_{\mathrm{CG}}\min\{n,\beta_{k}\log(\alpha_{\star}^{-1})\}+\sum_{k=1}^{\lfloor K_{2}\rfloor+1}C_{\mathrm{CG}}\min\{n,\beta_{k+k^{\star}}\log(\alpha_{\star}^{-1})\}$. Moreover, we call sqrt-cond at the zeroth iteration and the $k^{\star}$-th iteration. Hence, the total number of HVP computations called by sqrt-cond is at most $\Lambda_{3}$, by Lemma A.4 and Eq. (22). One HVP is computed in each iteration additionally. Putting these pieces together, the desired result follows from Eq. (21) and Eq. (22). ∎
We are ready to prove the convergence of IPPM, which shares a similar spirit with the methods from [Carmon-2018-Accelerated].
Proof of Theorem A.3.
With probability at least $1-(K+2)\delta^{\prime}\geq 1-\delta$, the results of Theorem A.10 hold for the first $\lfloor K\rfloor+2$ calls of Algorithm 3. Hence, if we are able to prove that IPPM must terminate in $\lfloor K\rfloor+2$ iterations under the assumption that the results of Theorem A.10 hold for all calls of Algorithm 3, then indeed IPPM terminates in $\lfloor K\rfloor+2$ iterations with probability at least $1-\delta$.

### v2

@@ A.3 Convergence Analysis
In this section, we prove the convergence of Algorithms 4 and 5. To do so, we first present some technical lemmas. Lemma A.4 analyzes the function sqrt-cond in Algorithm 4 by a classical result [38] on the Lanczos method. Lemma A.5 analyzes the function CG-Inverse in Algorithm 4 using [60, Theorem 5.5]. Lemma A.6 describes the relationship between the condition numbers of the Hessians at two nearby points for self-concordant functions.
@@
Proof Let $\epsilon^{\prime}=\frac{1}{10\beta^{2}}$. It follows from [38, Theorem 4.2] that, with probability at least $1-\frac{1}{2}\delta$, the Lanczos method outputs $\lambda_{1}$ such that $|\lambda_{1}-\lambda_{\max}(A)|\leq\frac{1}{3}\lambda_{\max}(A)$ in $O(\min\{n,\,\log(n\delta^{-2})\})$ matrix-vector multiplications. With probability at least $1-\frac{1}{2}\delta$, the Lanczos method outputs $\lambda^{\prime}$ such that $\left|\lambda^{\prime}-\lambda_{\max}\left(\tfrac{2\lambda_{1}}{1-\epsilon^{\prime}}I-A\right)\right|\leq\epsilon^{\prime}\cdot\allowbreak\lambda_{\max}\left(\tfrac{2\lambda_{1}}{1-\epsilon^{\prime}}I-A\right)\leq\tfrac{1}{3}\lambda_{\min}(A)$ in $O(\min\{n,\beta\log(n\delta^{-2})\})$ matrix-vector multiplications. Letting $\lambda_{2}=\frac{2\lambda_{1}}{1-\epsilon^{\prime}}-\lambda^{\prime}$, then we have $|\lambda_{2}-\lambda_{\min}(A)|\leq\frac{1}{3}\lambda_{\min}(A)$. Thus, letting $X^{2}=2\lambda_{1}/\lambda_{2}$ suffices. $\square$
Lemma A.5.
@@
Proof We know from [58, Theorem 5.1.13] that $x_{f}^{\star}$ exists. It follows from [58, Theorems 5.1.7 5.2.1] that $(1-\kappa r)^{2}\nabla^{2}f(x)\preceq\nabla^{2}f(x_{f}^{\star})\preceq\frac{1}{(1-\kappa r)^{2}}\nabla^{2}f(x)$, where $r=\|x-x_{f}^{\star}\|_{\nabla^{2}f(x)}\leq\frac{1}{(R^{-1}-1)\kappa}$. The desired result follows from direct calculation. $\square$
@@
Suppose $f$ is $\kappa$-self-concordant. At any given iteration of Algorithm 4, if $\beta_{k}^{2}\geq\mathrm{cond}(\nabla^{2}f(x_{k}))$ and $\rho_{k}>\frac{R_{1}^{2}}{\kappa^{2}}$, then $f(x_{k})-f(x_{k+1})\geq\kappa^{-2}\,\omega\left(\tfrac{\sqrt{1-\alpha_{\star}}}{1+\alpha_{\star}}\,\kappa\sqrt{\rho_{k}}\right)\geq\kappa^{-2}\omega(R_{3})$ and $\sqrt{\tfrac{\mathrm{cond}(\nabla^{2}f(x_{k+1}))}{\mathrm{cond}(\nabla^{2}f(x_{k}))}}\leq B_{k}\leq\exp\left(C_{1}\kappa^{2}(f(x_{k})-f(x_{k+1}))\right)$.
@@
Proof It follows from Lemma A.5 that ${\frac{\rho_{k}}{\delta_{k}}}\in[\frac{1-\alpha_{\star}}{(1+\alpha_{\star})^{2}},\frac{1+\alpha_{\star}}{(1-\alpha_{\star})^{2}}]$. Thus, Proposition 2.6 implies that $f(x_{k})-f(x_{k+1})\geq\kappa^{-2}\omega\left(\tfrac{\kappa\rho_{k}}{\sqrt{\delta_{k}}}\right)\geq\kappa^{-2}\omega\left(\tfrac{\sqrt{1-\alpha_{\star}}}{1+\alpha_{\star}}\kappa\sqrt{\rho_{k}}\right)\geq\kappa^{-2}\omega(R_{3})$. Let $r_{k}=\|x_{k+1}-x_{k}\|_{\nabla^{2}f(x_{k})}=\frac{\rho_{k}\sqrt{\delta_{k}}}{\delta_{k}+\kappa\rho_{k}\sqrt{\delta_{k}}}$. Hence, we have $(1-\kappa r_{k})^{-1}\leq 1+\frac{\sqrt{1+\alpha_{\star}}}{1-\alpha_{\star}}\kappa\sqrt{\rho_{k}}$. It follows from [58, Theorem 5.1.7] that $\sqrt{\tfrac{\mathrm{cond}(\nabla^{2}f(x_{k+1}))}{\mathrm{cond}(\nabla^{2}f(x_{k}))}}\leq(1-\kappa r_{k})^{-2}\leq B_{k}$. Moreover, we have $\tfrac{\log B_{k}}{\kappa^{2}(f(x_{k})-f(x_{k+1}))}\leq\tfrac{2\log\left(1+\frac{\sqrt{1+\alpha_{\star}}}{1-\alpha_{\star}}\,\kappa\sqrt{\rho_{k}}\right)}{\omega\left(\tfrac{\sqrt{1-\alpha_{\star}}}{1+\alpha_{\star}}\kappa\sqrt{\rho_{k}}\right)}\leq\tfrac{2\log\left(1+\frac{\sqrt{1+\alpha_{\star}}}{1-\alpha_{\star}}R_{1}\right)}{\omega\left(\tfrac{\sqrt{1-\alpha_{\star}}}{1+\alpha_{\star}}R_{1}\right)}\leq C_{1}$, where the second inequality holds by the monotonicity of $\frac{\log\left(1+\frac{\sqrt{1+\alpha_{\star}}}{1-\alpha_{\star}}t\right)}{\omega\left(\frac{\sqrt{1-\alpha_{\star}}}{1+\alpha_{\star}}t\right)}$ for $t>R_{1}$. This completes the proof. $\square$
@@
Proof If $\rho_{k}\leq\frac{R_{1}^{2}}{\kappa^{2}}$, then we have $t_{k}=\frac{\rho_{k}}{\delta_{k}+2\kappa\rho_{k}\sqrt{\delta_{k}}}$. Letting $\lambda_{k}=\lambda_{f}(x_{k})$, $H_{k}=\nabla^{2}f(x_{k})$, $G_{k}=\int_{0}^{1}\nabla^{2}f(x_{k}+\tau(x_{k+1}-x_{k}))d\tau$, $\zeta_{k}=\kappa\sqrt{\delta_{k}}$, we have $\nabla f(x_{k+1})=G_{k}(x_{k+1}-x_{k})+\nabla f(x_{k})=t_{k}G_{k}h_{k}+g_{k}$. It follows from [58, Corollary 5.1.5] that $(1-t_{k}\zeta_{k}+\frac{1}{3}t_{k}^{2}\zeta_{k}^{2})H_{k}\preceq G_{k}\preceq\frac{1}{1-t_{k}\zeta_{k}}H_{k}$. Moreover, [58, Lemma 5.1.7] implies that $H_{k+1}^{-1}\preceq\frac{1}{(1-t_{k}\zeta_{k})^{2}}H_{k}^{-1}$. Thus, we have $\lambda_{f}(x_{k+1})^{2}\leq\tfrac{1}{(1-t_{k}\zeta_{k})^{2}}\|\nabla f(x_{k+1})\|_{H_{k}^{-1}}^{2}$, where
$\|\nabla f(x_{k+1})\|_{H_{k}^{-1}}^{2}=\lambda_{k}^{2}-2t_{k}h_{k}^{\top}G_{k}h_{k}+t_{k}^{2}h_{k}^{\top}G_{k}H_{k}^{-1}G_{k}h_{k}+2t_{k}h_{k}^{\top}G_{k}(H_{k}^{-1}g_{k}+h_{k}).$
@@
$|2h_{k}^{\top}G_{k}(H_{k}^{-1}g_{k}+h_{k})|\leq\alpha_{\star}\|G_{k}h_{k}\|_{H_{k}^{-1}}^{2}+\alpha_{\star}^{-1}\|H_{k}^{-1}g_{k}+h_{k}\|_{H_{k}}^{2}.$
@@
Lemma A.5 implies $\zeta_{k}\in[0,\frac{1+\alpha_{\star}}{\sqrt{1-\alpha_{\star}}}R_{1}]$. For any $\zeta_{l}<\zeta_{u}$, we define $t_{u}=\frac{1}{\frac{(1-\alpha_{\star})^{2}}{1+\alpha_{\star}}+3\zeta_{l}}$ and $t_{l}=\frac{1}{\frac{(1+\alpha_{\star})^{2}}{1-\alpha_{\star}}+3\zeta_{u}}$. Since $t_{k}=\frac{1}{\frac{\delta_{k}}{\rho_{k}}+3\zeta_{k}}$ and $\frac{\delta_{k}}{\rho_{k}}\in[\frac{(1-\alpha_{\star})^{2}}{1+\alpha_{\star}},\frac{(1+\alpha_{\star})^{2}}{1-\alpha_{\star}}]$, we have $g(t_{k},\zeta_{k})\leq\tfrac{1-2t_{l}(1-\alpha_{\star})^{2}(1-t_{u}\zeta_{u})}{(1-t_{l}\zeta_{l})^{2}}+\tfrac{t_{u}^{2}(1+\alpha_{\star})^{2}}{(1-t_{u}\zeta_{u})^{4}}+\tfrac{\alpha_{\star}t_{u}}{{(1-t_{u}\zeta_{u})^{2}}}\left(\tfrac{(1+\alpha_{\star})^{2}}{(1-t_{u}\zeta_{u})^{2}}+1\right)$ for any $\zeta_{k}\in[\zeta_{l},\zeta_{u}]$. Dividing $0,\frac{1+\alpha_{\star}}{\sqrt{1-\alpha_{\star}}}R_{1}]$ into $1000$ intervals equally, we can verify numerically that $g(t_{k},\zeta_{k})\leq 0.945$. The desired result follows from Lemma A.5. $\square$
@@
Lemma A.9 shows that $A_{1}\cap A_{2}\cap A_{3}$ happens with high probability. Hence, using two calls of sqrt-cond, Algorithm 4 gets good estimates of Hessian condition numbers. If $A_{1}\cap A_{2}\cap A_{3}$ holds, then for all $0\leq k\leq k^{\star}$, using Lemma A.7, we have
@@
Proof By Lemma A.4, we have $\mathbb{P}\left(A_{1}\right)\geq 1-\frac{1}{2}\delta$. Therefore, it suffices to show $\mathbb{P}\left(A_{2}|A_{1}\right)\geq 1-\frac{1}{2}\delta$ and $A_{1}\cap A_{2}\subseteq A_{3}$. To show $\mathbb{P}\left(A_{2}|A_{1}\right)\geq 1-\frac{1}{2}\delta$, we observe from Lemma A.7 that $\beta_{0}^{2}\geq\mathrm{cond}(\nabla^{2}f(x_{0}))$ implies $\beta_{k^{\star}}^{2}\geq\mathrm{cond}(\nabla^{2}f(x_{k^{\star}}))$. Thus, Lemma A.4 implies $\mathbb{P}\left(A_{2}|A_{1}\right)\geq 1-\frac{1}{2}\delta$. It remains to show $A_{1}\cap A_{2}\subseteq A_{3}$. Suppose that $A_{1}\cap A_{2}$ holds, then we have $\mathrm{cond}(\nabla^{2}f(x_{l}))\leq\beta_{l}^{2}$ for $0\leq l\leq k^{\star}$ by Lemma A.7. Moreover, Lemma A.8 yields $\rho_{l}\leq\frac{R_{1}^{2}}{\kappa^{2}}$ for all $l\geq k^{\star}+1$. Therefore, using Lemma A.6, we have $\mathrm{cond}(\nabla^{2}f(x_{l}))\leq C_{3}^{8}\mathrm{cond}(\nabla^{2}f(x_{k^{\star}}))\leq C_{3}^{8}(\beta^{\star})^{2}=\beta_{\mathrm{local}}^{2}=\beta_{l}^{2}$. $\square$
@@
Suppose that $f$ is $\kappa$-self-concordant, bounded below, and that $\beta^{2}\geq\mathrm{cond}(\nabla^{2}f(x_{0}))$. Then, with probability at least $1-\delta$, the following statements hold: 1. Algorithm 4 terminates and returns a point $y$. If $y\neq x_{0}$, then $\lambda_{f}(y)\leq\epsilon_{1}$ and $\lambda_{f}(x_{0})\geq\sqrt{\frac{1-\alpha_{\star}}{1+\alpha_{\star}}}\,\epsilon_{0}$; if $y=x_{0}$, then $\lambda_{f}(x_{0})\leq\epsilon_{0}$. Let $\Delta_{f}=\kappa^{2}(f(x_{0})-f(y))$, $K_{1}=\frac{\Delta_{f}}{\omega(R_{3})}$, and $K_{2}=2\log_{C_{2}^{-1}}\!\left(\frac{R_{2}}{\kappa\epsilon_{1}}\right)$. Then Algorithm 4 terminates within $2+\lfloor K_{1}\rfloor+\max(\lfloor K_{2}\rfloor,0)$ iterations. 2. Let $B=\sqrt{\max\left\{{\mathrm{cond}(\nabla^{2}f(x_{0}))},\mathrm{cond}(\nabla^{2}f(\mathop{\rm argmin}f))\right\}}$, $\Lambda_{1}=C_{\mathrm{CG}}\min\left\{n,2Be^{C_{1}\Delta_{f}}\log(\alpha_{\star}^{-1})\right\}$, $\Lambda_{2}=C_{\mathrm{CG}}\min\left\{n,2BC_{3}^{6}\log(\alpha_{\star}^{-1})\right\}$, and $\Lambda_{3}=C_{L}\min\left\{n,\beta\log(n\delta^{-2})\right\}+C_{L}\min\left\{n,2Be^{C_{1}\Delta_{f}}\log(n\delta^{-2})\right\}$. The total number of HVP computations performed by Algorithm 4 is at most $(\Lambda_{1}+1)(K_{1}+1)+(\Lambda_{2}+1)(K_{2}+1)+\Lambda_{3}=O(\log(\epsilon_{1}^{-1})+\log(\delta^{-1}))$.
@@
Proof By Lemma A.9, it suffices to prove all statements hold deterministically assuming that the event $A_{1}\cap A_{2}\cap A_{3}$ holds. For all $k=0,\ldots,k^{\star}-1$, we have $\rho_{k}>\frac{R_{1}^{2}}{\kappa^{2}}$. Therefore, it follows from Lemma A.7 that $f(x_{k})-f(x_{k+1})\geq\kappa^{-2}\omega(R_{3})$. Summing this inequality over $k=0,1,\dots,k^{\star}-1$ yields ${k^{\star}}\kappa^{-2}\omega(R_{3})\leq f(x_{0})-f(x_{k^{\star}})$. Thus, $k^{\star}$ must be finite. Lemma A.8 implies that we have $\rho_{k+1}\leq C_{2}\rho_{k}\leq\frac{R_{1}^{2}}{\kappa^{2}}$ for all $k\geq k^{\star}$. Thus, we have $(1-\alpha_{\star})\epsilon_{1}^{2}\leq\rho_{k}\leq C_{2}^{k-k^{\star}}\rho_{k^{\star}}\leq C_{2}^{k-k^{\star}}\frac{R_{1}^{2}}{\kappa^{2}}$. Therefore, the algorithm terminates at some $k\leq 1+\lfloor K_{1}\rfloor+\max(\lfloor K_{2}\rfloor,0)$.
@@
Now, we can estimate the total number of HVPs. By Lemma A.5, the total number of HVPs called by CG-Inverse is bounded by $\sum_{k=0}^{k^{\star}}C_{\mathrm{CG}}\min\{n,\beta_{k}\log(\alpha_{\star}^{-1})\}+\sum_{k=1}^{\lfloor K_{2}\rfloor+1}C_{\mathrm{CG}}\min\{n,\beta_{k+k^{\star}}\log(\alpha_{\star}^{-1})\}$. Moreover, we call sqrt-cond at the zeroth iteration and the $k^{\star}$-th iteration. Hence, the total number of HVP computations called by sqrt-cond is at most $\Lambda_{3}$, by Lemma A.4 and Eq. (19). One HVP is computed in each iteration additionally. Putting these pieces together, the desired result follows from Eq. (18) and Eq. (19). $\square$
We are ready to prove the convergence of IPPM, which shares a similar spirit with [9].
Proof.
Proof of Theorem A.3 With probability at least $1-(K+2)\delta^{\prime}\geq 1-\delta$, the results of Theorem A.10 hold for the first $\lfloor K\rfloor+2$ calls of Algorithm 4. Hence, if we are able to prove that IPPM must terminate in $\lfloor K\rfloor+2$ iterations under the assumption that the results of Theorem A.10 hold for all calls of Algorithm 4, then indeed IPPM terminates in $\lfloor K\rfloor+2$ iterations with probability at least $1-\delta$.